# White Men Lead, Black Women Help? Benchmarking Language Agency Social Biases in LLMs

### v4

@@ D.2 Stratified Analysis on Human-Written Datasets
D.3 Full Evaluation Results
@@
Table  17 shows results for gender biases before mitigation, whereas Table  18 presents results after mitigation. Table  19 presents results for racial biases before mitigation, and Table  20 shows results after mitigation. For intersectional biases, results for ChatGPT before mitigation are in Table  21; results after mitigation are in Table  22. Intersectional results for Mistral before mitigation are in Table  23; results after mitigation are in Table  24. Intersectional outcomes for Llama3 before mitigation are in Table  25; results after mitigation are in Table  26.
Appendix E Computational Resources
@@
Reference Letter (Wan et al., 2023) Male 57.47 42.53 14.94 4.64

### v5

@@ E.2.2 Stratified Analysis on Human-Written Datasets
E.2.3 Full Evaluation Results
@@
Table  19 shows results for gender biases before mitigation, whereas Table  20 presents results after mitigation. Table  21 presents results for racial biases before mitigation, and Table  22 shows results after mitigation. For intersectional biases, results for ChatGPT before mitigation are in Table  23; results after mitigation are in Table  24. Intersectional results for Mistral before mitigation are in Table  25; results after mitigation are in Table  26. Intersectional outcomes for Llama3 before mitigation are in Table  27; results after mitigation are in Table  28.
Appendix F Computational Resources
@@
Reference Letter (Wan et al., 2023a) Male 57.47 42.53 14.94 4.64